Solve the equation
1/4(8x - 12) = 6[1/3x - 1/2] + 10
I have no idea on how to even go about solving this. Please help.

Answer by wuwei96815(245)

1/4(8x-12)=6[1/3x-1/2]+10
First remove the parentheses and brackets by multiplying everything within by what is outside of them, with the exception of the 10 which is a separate term.
2x-3=2x-3+10
Now add all like terms.
2x-2x=3-3+10
0=10
The answer does not make any sense. What you started with was not a true equation. Where did you get this problem?